# Are Vision Language Models Texture or Shape Biased and Can We Steer Them?

### v1

@@ Abstract
Vision language models (VLMs) have drastically changed the computer vision model landscape in only a few years, opening an exciting array of new applications from zero-shot image classification, over to image captioning, and visual question answering. Unlike pure vision models, they offer an intuitive way to access visual content through language prompting. The wide applicability of such models encourages us to ask whether they also align with human vision — specifically, how far they adopt human-induced visual biases through multimodal fusion, or whether they simply inherit biases from pure vision models. One important visual bias is the texture vs. shape bias, or the dominance of local over global information. In this paper, we study this bias in a wide range of popular VLMs. Interestingly, we find that VLMs are often more shape-biased than their vision encoders, indicating that visual biases are modulated to some extent through text in multimodal models. If text does indeed influence visual biases, this suggests that we may be able to steer visual biases not just through visual input but also through language: a hypothesis that we confirm through extensive experiments. For instance, we are able to steer shape bias from as low as 49% to as high as 72% through prompting alone. For now, the strong human bias towards shape (96%) remains out of reach for all tested VLMs.
Figure 1: Unlike many unimodal models, vision language models (VLMs) prefer shape over texture for object recognition, but not to the same extent as humans. Further, we find that the (visual) texture/shape bias [1] can be steered through language alone, albeit not to the extent as through vision. Here we visualize the texture/shape bias of some exemplary VLMs, and highlight the steerability of InternVL-Chat 1.1 [2].
@@ 1 Introduction
As the old adage goes, all models are wrong, but some are useful. Similarly, all (machine learning) models are biased, but according to the no free lunch theorem, only some biases are useful [3, 4]. While most biases observed in language models reflect social particularities that originate in unbalanced training data, vision models have been shown to be exceptionally misaligned with human perception, referred to as the texture vs. shape bias [1]. Specifically, when recognizing objects, an object’s shape often plays a minor role to models while the texture is strongly preferred. Humans instead predominantly decide by the object’s shape ($96\%$ shape over texture decisions). This finding has received significant attention [5, 6, 7, 8, 9, 10], because it shows that our vision models, while approaching near-human levels in accuracy on specialized tasks, poorly approximate fundamentals of human perception.
As the current generation of deep learning models is increasingly multimodal [11, 12, 13, 14], it is often unclear whether these models inherit certain biases from their text encoders, vision encoders, or whether those interact and are combined through multimodal fusion. Specifically for the (vision-only) texture vs. shape bias, we are interested in how this bias is influenced by language. If language is indeed able to influence a vision-only bias, this offers the possibility of aligning and generally steering biases simply and intuitively via prompts.
We study of the texture vs. shape bias and its steerability in recent vision language models (VLMs). Here we use VLM to refer to text/vision-input models that output text generated by a large language model (LLM). Our investigation shows that the model’s inherent bias towards texture is far less pronounced in VLMs than in most previously studied vision-only models. As shown in Fig. 1, most VLMs decide by shape more often than by texture, even when prompted in a neutral way (while not approaching the human shape bias). Further, we can confirm that VLMs learn to understand the visual concepts of shape and texture, such that they allow steering of the prediction to some extent by simple prompt modifications.
Our contribution can be summarized into the following findings:
  • Text appears to influence visual texture/shape bias. Through a large-scale study, we report that even though texture bias is a vision-only bias found in common vision models, VLMs are often more shape-biased than their vision encoders, even though they decide based on texture more often than humans would do (Sec. 4).
  • VLMs have an understanding of the visual concepts shape and texture which allows steering. Given that multimodal models do not appear to simply inherit visual biases from their vision encoders, this suggests that we may be able to steer visual biases through language. We test and confirm this hypothesis with a set of experiments. For instance, shape bias can be steered as low as 49% and as high as 72% through prompting alone: given that visual biases can be influenced through text, this insight can be harnessed to steer them (Sec. 5).
@@ 2 Related Work
Vision Language Models. Sparked by the success of vision language pretraining [11, 15, 16, 17], where features extracted from image-text paired data are aligned in a joint embedding space, recent VLMs added language modeling during training [18, 19, 20], enabling models to reason about images. Subsequently, finetuning these VLMs on instruction-following data [13, 21, 22, 23, 14], such as reinforcement learning from human feedback (RLHF) [24], enables users to prompt these models, easing their usability for humans. Resulting models are commercialized [12, 25, 26] or open-sourced [21, 23, 2], and consequently become accessible to a wide range of users.
@@
Shape Bias. Humans primarily rely on shape information to recognize objects. On the other hand standard ImageNet-trained vision-models, such as convolutional neural networks (CNNs), are biased towards texture to make their classification decisions [1]. Nonetheless, shape information can still be present in layers/latent space of the model before the classifier [5, 7]. Prior research has shown that the texture bias of CNNs can be reduced in training [1, 27, 28, 5, 29, 30, 31]. But the network architecture has a high influence, and vision-only ViTs [32] were shown to be more shape-biased by default [9], more human-like [33], scalable by data size [34], and can be explicitly designed to separate shape and texture in their token space [9]. Jointly embedding vision and language in these networks (but not CNNs as we will show!), through CLIP [11], further increases their shape bias in zero-shot classification [29]. Yet, these models still do not reach human levels. The only known models to achieve such levels are image-generative classifiers [31], which in some way also combine vision and language.
Other Biases. For the sake of completeness, we list other relevant works on uni/multimodal biases. However, we want to point out that the focus of this paper lies on the texture/shape bias.
Vision biases have been widely investigated, such as single-demographic effects (race and gender) for face recognition tasks [35, 36], or high-frequency biases for classification tasks [37, 27]. Recently, [10] found another bias that separates human from model vision, by investigating the critical frequency band of object recognition. For language models, several works focus on investigating societal biases, such as gender and race [38, 39] and ways of debiasing them [39, 40, 41], or explicitly forcing them [42]. A recent study also found that LLMs are biased towards high-value over likely options [43]. Another study focused on encoded moral beliefs [44]. LLMs can also pick up human traits - one study found that adding "take a deep breath" to prompts improves performance [45]. Of course, some of the unimodal biases also apply to VLMs (e.g., [45]), but a few works have also explicitly focused on biases in VLMs. For example, neurons of CLIP [11] were studied in [46], revealing that some neurons respond to the same concept regardless of its presentation, which is a potential reason for the high generalizability. On the other hand, this enables attacks by rendering text on images (typographic attacks). Additionally, several works demonstrated that VLMs fail to count objects [11, 47, 48], and generally struggle in structured tasks [34].
3 Methodology
@@
A cornerstone of our analysis is the measurement of texture/shape bias in (LLM-based) VLMs when performing tasks that require some level of image classification. In the following subsections, we introduce how and in which tasks exactly we measure shape bias.
3.1 Measuring Texture/Shape Bias
@@
We use the texture-shape cue-conflict classification problem (cue-conflict) [1] consisting of 1,280 samples with conflicting shape and texture cues synthetically generated via a style transfer model [49] from ImageNet [50] samples (see Fig. 1 for examples). The shape and texture classes belong to 16 super-classes of ImageNet. Following [1], we have excluded 80 images from the dataset where texture and shape cues belong to the same class.
We want to emphasize that from an information perspective alone, predicting either label (or both) would be correct. However, humans tend to prioritize the shape cue for predictions which is in stark contrast to most models [1].
@@
Throughout the paper, we will refer to this as the accuracy. We use the definition of shape bias [1], which is defined by the ratio of shape decisions over accurate decisions:
Shape Bias=Shape $\textit{Shape Bias}=\frac{\textit{Shape~{}Accuracy}}{Cue~{}Accuracy}$ (2)
While we primarily focus on measuring the shape bias in this study, the accuracy provides an important signal of the model quality and robustness, as well as for the comparability of results.
3.2 VLM Tasks
We measure the texture/shape bias of VLMs in two common tasks: visual question answering (VQA) [51] where we seek to obtain a zero-shot classification [11] of the object, and image captioning [52] where we look for an accurate description of the image. In both cases, we use the cue-conflict dataset in single-round answering with no shared conversation history between conversations.
3.2.1 Image Captioning.
@@ 3.2 VLM Tasks
As the responses are open-ended we rely on zero-shot classifications of the generated description to marginalize out the most descriptive class. To this end, we embed the generated descriptions and all (raw) class labels using ember-v1 [53] and predict the class with the smallest cosine distance (similar to zero-shot classification in CLIP [11]). However, the generated caption may refer to multiple class labels (or none). As an additional signal, we perform a more granular analysis using an additional LLM (Nous-Hermes-2-Mixtral-8x7B-DPO [54]) by instructing the model to extract all mentioned classes (similar to [55]). This allows us to understand if the model detects both cues (but the embedding model enforces a specific prediction) and to quantify how often the model response is too generic to detect any class.
3.2.2 VQA Classification.
Following the questioning style in LLaVA [56], we ask the model "Which option best describes the image?" and provide an alphabetic enumeration of all class labels in the style "A. airplane". For a simpler response extraction and confidence evaluation (see below), we end the prompt by instructing the model to answer with only the letter corresponding to the correct answer ("Answer with the option’s letter from the given choices directly."). Compared to captioning, this is similar to the discrimination in ImageNet [50] image classifiers [57, 58, 59, 60, 32] in the sense that it only allows the model to respond with a single class and does not provide an option to not answer - if models follow the instruction.
Response Extraction.
Despite instructing the models to only respond with an option letter, we observe multiple response styles: option letter + label ("H. cat."), just the label ("cat."), long explanation containing the option letter and/or label ("The image features a black and white image of a cat."). In all cases, punctuation and capitalization may be different ("H.", "H", "h)"). The first two response styles are easily correctable by simple post-processing (we prioritize the option letter in case of a conflicting option letter and label), and in some cases, explanations can be corrected as well if the response includes the option letter. However, we avoid heavy post-processing and consider individual answers as wrong if they are not recoverable. In most cases, the ratio of these is negligible.
Confidence of VQA predictions.
All answer options in our prompts correspond to a single character and, thus, token. Well-behaving models, where the response consistently starts with the option letter (and nothing else) allow us to gather insights into the prediction process. For these models, the logits of each token correspond to the logits of option letters. By applying a softmax function we can then analyze the confidence in each option (we map invalid tokens to a separate null class). This allows us to perform additional studies on sampling behavior.
4 Are VLMs Biased towards Texture or Shape?
In this section, we measure the shape bias of VLMs and try to understand if there are driving factors (specific models, scale, RLHF [24] etc.) behind potential differences in behavior. Beyond this quantitative analysis, we aim to better understand the internal mechanisms through insights into the decision formation and ablation of the representation of the vision encoder.
4.1 Main Results
We start our experimental evaluation by measuring the shape bias in the VQA and captioning tasks, using a collection of diverse vision language models reflecting the multitude of research directions. These models include connections of common pretrained CLIP encoders and LLMs [21, 61, 62, 23, 63, 64, 65, 66], mixture-of-expert-LLMs [67], optimized architectures for resource-constrained systems [68], finetuning with RLHF [69, 12], massive vision encoders [2]. Additionally, we survey commercial, closed-source models like Gemini Pro Vision 1.0 [25], GPT-4V (Preview) [12], and Qwen-VL Plus/Max [26] where access is limited to APIs and few details are known. For a detailed list of models, we refer the reader to the appendix.
@@ 4.1 Main Results
Figure 2: Most VLMs are slightly shape-biased but some models show differences when asked to describe an image compared to VQA. We measure the shape bias on the cue-conflict dataset [1]. For reference, we also provide measurements on ResNet-50 [59] from the initial shape bias study [1], zero-shot classification (CLIP ViT-L/14 [11]), and a human average (over 10 subjects [1]).

### v2

# Can We Talk Models Into Seeing the World Differently?

@@ Abstract
Unlike traditional vision-only models, vision language models (VLMs) offer an intuitive way to access visual content through language prompting by combining a large language model (LLM) with a vision encoder. However, both the LLM and the vision encoder come with their own set of biases, cue preferences, and shortcuts, which have been rigorously studied in uni-modal models. A timely question is how such (potentially misaligned) biases and cue preferences behave under multi-modal fusion in VLMs. As a first step towards a better understanding, we investigate a particularly well-studied vision-only bias - the texture vs. shape bias and the dominance of local over global information. As expected, we find that VLMs inherit this bias to some extent from their vision encoders. Surprisingly, the multi-modality alone proves to have important effects on the model behavior, i.e., the joint training and the language querying change the way visual cues are processed. While this direct impact of language-informed training on a model’s visual perception is intriguing, it raises further questions on our ability to actively steer a model’s output so that its prediction is based on particular visual cues of the user’s choice. Interestingly, VLMs have an inherent tendency to recognize objects based on shape information, which is different from what a plain vision encoder would do. Further active steering towards shape-based classifications through language prompts is however limited. In contrast, active VLM steering towards texture-based decisions through simple natural language prompts is often more successful.
URL: https://github.com/paulgavrikov/vlm_shapebias
Figure 1: Language can be used to steer visual cue preferences (biases) in vision language models (VLMs). Here we illustrate the (visual) texture/shape bias Geirhos et al. (2019) of some exemplary VLMs, and highlight the steerability of InternVL-Chat 1.1 Chen et al. (2024) through the processing of vision and text inputs (prompts).
@@ 1 Introduction
As the old adage goes, all models are wrong, but some are useful. Similarly, recent machine learning models have proven to be very useful in practice, although we know their decisions to be impacted by specific biases, such as cue preferences misaligned with human perception and shortcuts Geirhos et al. (2020a). Some of these cue biases are particularly misaligned in traditional, uni-modal models and often reveal fundamental differences in the decision function compared to humans Geirhos et al. (2019); Subramanian et al. (2024); Wang et al. (2020); Buolamwini & Gebru (2018); Raji & Buolamwini (2019). However, the current generation of deep learning models is increasingly multi-modal, for example, by fusion of large language models (LLMs) with modality-specific encoders OpenAI (2023); Alayrac et al. (2022); Huang et al. (2023). On the one hand, this approach allows for an exciting array of applications that can be defined at inference via prompts in natural language. On the other hand, the once well-studied biases are now combined in multi-modal fusion, leaving open questions on how and if the specific biases interact.
Specifically for vision language models (VLMs) we are therefore asking if we can talk models into seeing the world differently - i.e., to what extent does the LLM-based multi-modal fusion change the cues predominantly used for image classification and, furthermore, can we utilize natural language prompts to override the inductive biases of vision encoders. If language is indeed able to influence a vision-only bias, this may offer the possibility of aligning model behavior (with human behavior) using intuitive language prompting.
In general, measuring biases in the visual cues used by a model to make a particular prediction is hard. While we assume that there is a multitude of cue biases learned in vision models, only a few of them are harmful or misaligned [^1] We are explicitly focusing on biases in terms of low-level vision cues such as shape versus texture or high frequency versus low frequency. High-level biases, that may have a societal impact, are therefore explicitly excluded from this investigation. Please see Appendix O for a discussion. . An example of a benign bias would be the "foreground" bias, i.e., models mostly classify images by their foreground objects. Similarly, objects in the image center are usually perceived to be more important than objects in the periphery. These biases follow human intuition and have therefore not been causing much controversial discussion. This is in contrast to the texture vs. shape bias Geirhos et al. (2019) - one of the best-studied cue biases in object recognition models Hermann et al. (2020); Shi et al. (2020); Islam et al. (2021); Benarous et al. (2023); Naseer et al. (2021); Subramanian et al. (2024). It states that humans predominantly recognize objects in images by their shape ($96\%$ shape over texture decisions), whereas vision models strongly prioritize texture cues and discount the object’s shape often. Machine perception is thus at odds with human intuition, even if the model accuracy is high.
We are the first to provide a large-scale study for VLMs, investigating the texture/shape bias in object recognition in visual question answering and image captioning. Our investigation shows that the texture bias is by default far less pronounced in VLMs than in most previously studied vision-only models. As shown in Fig. 1, VLMs decide by shape more often than by texture - albeit not matching the human shape bias ($96\%$). Further, we find that by using biased instructions, we can steer the model output to some extent in both directions, toward a texture or shape bias. This demonstrates that visual biases in multi-modal models can be influenced by language, opening up an exciting new possibility of aligning model outputs using language prompts, without the need for retraining. In summary, our large-scale study offers the following findings:
  • We show that VLMs preserve cue biases of their vision-encoders only to some extent, e.g., they are yielding decisions that are more shape-biased than pure vision models, albeit not reaching human levels of the shape bias (Sec. 4.1).
  • We show that uni-modal vision encoders generate a flexible representation that contains texture and shape cues. Ultimately, the language modality (through the LLM) tends to suppress either of the cues, such that objects are recognized purely on the grounds of either shape or texture (Sec. 4.2).
  • We find that VLMs offer a unique opportunity to steer visual biases through language alone, stemming from the multi-modal fusion. For instance, we can steer the shape bias as low as 49% and as high as 72% through prompting alone without significantly affecting the accuracy (Sec. 5). This ability is not limited to the texture/shape bias (Sec. 5.2).
@@ 2 Related Work
Sparked by the success of vision language pretraining Radford et al. (2021); Jia et al. (2021); Zhai et al. (2022b); Sun et al. (2023b), where features extracted from image-text paired data are aligned in a joint embedding space, recent VLMs added language modeling during training Li et al. (2022); Yu et al. (2022); Li et al. (2023b), enabling models to reason about images. Subsequently, finetuning these VLMs on instruction-following data Alayrac et al. (2022); Liu et al. (2023b); Luo et al. (2023); Dai et al. (2023); Huang et al. (2023), such as reinforcement learning from human feedback (RLHF) Ouyang et al. (2022), enables users to prompt these models, easing their usability for humans. Resulting models are commercialized OpenAI (2023); Gemini Team (2023); Qwen Team (2024) or open-sourced Liu et al. (2023b); Dai et al. (2023); Chen et al. (2024), and consequently become accessible to a wide range of users.
@@
This success of vision language models calls for improving our understanding of the visual cues leveraged by VLMs and the degree to which these can be affected through language. In particular, the fact that vision models leverage cues different from the ones intuitively used by humans has been widely discussed. In our study, we focus on the texture vs. shape bias, as a particularly well-studied example in vision-only models. Humans primarily rely on shape information to recognize objects. This is in contrast to standard ImageNet-trained vision models, such as convolutional neural networks (CNNs), which are biased towards texture to make their classification decisions Geirhos et al. (2019). Nonetheless, shape information can still be present in layers/latent space of the model before the classifier Hermann et al. (2020); Islam et al. (2021). Prior research has shown that the texture bias of CNNs can be reduced in training Geirhos et al. (2019); Lukasik et al. (2023); Li et al. (2021); Hermann et al. (2020); Geirhos et al. (2021); Gavrikov et al. (2023); Jaini et al. (2024); Gavrikov & Keuper (2024). But the network architecture has a high influence, and vision-only ViTs Dosovitskiy et al. (2021) were shown to be more shape-biased by default Naseer et al. (2021), more human-like Tuli et al. (2021), scalable by data size Zhai et al. (2022a), and can be explicitly designed to separate shape and texture in their token space Naseer et al. (2021). Jointly embedding vision and language in these networks (but not CNNs), through CLIP Radford et al. (2021), further increases their shape bias in zero-shot classification Geirhos et al. (2021). Yet, these models still do not reach human levels. The only known models to achieve such levels are image-generative classifiers Jaini et al. (2024), which also combine vision and language in a different manner.
Measuring Texture/Shape Bias.
@@
A cornerstone of our analysis is the measurement of the texture/shape bias in (LLM-based) VLMs when performing tasks that are based on object recognition. In the following, we summarize how this bias is measured for vision-only models, which forms the basis for our study.
@@
Like most studies on the texture/shape bias in vision models, we use the texture-shape cue-conflict classification problem (cue-conflict) Geirhos et al. (2019) consisting of 1,280 samples with conflicting shape and texture cues synthetically generated via a style transfer model Gatys et al. (2016) from ImageNet Deng et al. (2009) samples (see Fig. 1 for examples). The shape and texture classes belong to 16 super-classes of ImageNet. Following Geirhos et al. (2019), we have excluded 80 images from the dataset where texture and shape cues belong to the same class. From an information perspective alone, predicting either label (or both) would be correct. However, humans tend to prioritize the shape cue for predictions, which is in stark contrast to most models Geirhos et al. (2019; 2021).
@@
Throughout the paper, we will refer to this as the accuracy. We use the definition of shape bias Geirhos et al. (2019), which is defined by the ratio of shape decisions over accurate decisions:
[…]
While we primarily focus on measuring the shape bias in this study, accuracy is an important signal for steering in later sections, as it demonstrates how the bias was obtained. For instance, achieving perfect shape bias in any model would be possible by simply mislabeling all texture-biased detections – but it would affect accuracy. Importantly, changes in accuracy (positive or negative) observed on the cue-conflict dataset may not necessarily generalize to other datasets.
3 Measuring Cue Biases in VLMs
Given a dataset such as proposed in Geirhos et al. (2019) for shapes and textures, we propose to measure the cue bias of VLMs in two tasks: visual question answering (VQA) Antol et al. (2015), where we seek to obtain a zero-shot classification Radford et al. (2021) of the object, and image captioning Vinyals et al. (2015) where we look for an accurate but brief description of objects in the image. For both tasks, we evaluate single-round answering with no shared conversation history between conversations.
3.1 VQA Classification
Following the questioning style in LLaVA Li et al. (2023a), we ask the model "Which option best describes the image?" and provide an alphabetic enumeration of all class labels in the style "A. airplane". For a simpler response extraction and confidence evaluation (see below), we end the prompt by instructing the model to answer with only the letter corresponding to the correct answer ("Answer with the option’s letter from the given choices directly."). Compared to captioning, this is similar to the discrimination in ImageNet Deng et al. (2009) image classifiers Krizhevsky et al. (2012); Srivastava et al. (2015); He et al. (2015); Huang et al. (2017); Dosovitskiy et al. (2021) in the sense that it only allows the model to respond with a single class and does not provide an option to not answer - if models follow the instruction.
Response Extraction. Despite instructing the models to only respond with an option letter, we observe multiple response styles: option letter + label ("H. cat."), just the label ("cat."), long explanation containing the option letter and/or label ("The image features a black and white image of a cat."). In all cases, punctuation and capitalization may be different ("H.", "H", "h)"). The first two response styles are easily correctable by simple post-processing (we prioritize the option letter in case of a conflicting option letter and label), and in some cases, explanations can be corrected as well if the response includes the option letter. However, we avoid heavy post-processing and consider individual answers wrong if they are not recoverable. In most cases, the ratio of these is negligible.
3.1.1 Active steering through Prompts
The above setting allows us to test the inherent cue bias of a given VLM. Yet, the multi-modal nature of VLMs paves the way to not only test for a given bias but also to actively steer the model towards using particular types of visual cues. Note that models are not trained to perform well under this task, and it is unclear how flexible they are in basing a particular decision on one or another type of cue (for example on texture or shape).
To explore the flexibility of model predictions for given types of visual cues, we conduct experiments comparing performance under a default neutral prompt and biased instructions.
In the simplest case, we can test the cue bias steering through hand-crafted prompting, where a model is asked to identify the class using a particular visual cue (e.g.,"Identify the primary shape in the image."). Details on our tested biased prompts can be found in Sec. C.3.
Automated Prompt Engineering. To further enhance the steering signal provided by the language prompt, we further evaluate automatically crafted prompts. This is achieved by employing an LLM as optimizer Yang et al. (2024) to continuously generate new prompts in natural language targeting to maximize either shape or texture bias in a feedback loop. We provide the LLM feedback about the achieved accuracy and shape bias. Additionally, we opt for greedy token sampling in the VLM to reduce noise in the feedback loop. For further details, we refer the reader to Appendix E.
3.2 Image Captioning
@@
Response Extraction. As the responses are open-ended, we rely on zero-shot classifications of the generated description to marginalize the most descriptive class. To this end, we embed the generated descriptions and all (raw) class labels using ember-v1 Nur & Aliyev (2024) and predict the class with the smallest cosine distance (similar to zero-shot classification in CLIP Radford et al. (2021)). However, the generated caption may refer to multiple class labels (or none). As an additional signal, we perform a more granular analysis using an additional LLM (Nous-Hermes-2-Mixtral-8x7B-DPO Teknium et al. (2024)) by instructing the model to extract all mentioned classes (similar to Yan et al. (2021)). This allows us to understand if the model detects both cues (but the embedding model enforces a specific prediction) and to quantify how often the model response is too generic to detect any class.
4 Cue Biases in VLMs: An Analysis of Texture versus Shape
@@
Figure 2: Most VLMs prioritize shapes over texture cues. We measure the shape bias on the cue-conflict dataset Geirhos et al. (2019). For reference, we also provide measurements on an ImageNet-trained ResNet-50 He et al. (2015), zero-shot classification with CLIP ViT-L/14 Radford et al. (2021), and a human average (over 10 subjects Geirhos et al. (2019)). The results in table format are shown in Appendix B.
We base our core analysis on the texture/shape cue-conflict dataset Geirhos et al. (2019) evaluation of the texture versus shape bias. Traditional image classification models like ResNet-50 trained on ImageNet have been shown to severely prioritize texture cues (only $22\%$ shape bias) starkly contrasting the strong shape bias of humans ($96\%$). As of now, results on multi-modal LLM-based models are missing. Therefore, we start our experimental evaluation by measuring the shape bias for the VQA and Image Captioning tasks, using a collection of diverse VLMs reflecting the multitude of research directions. These models include connections of common pretrained CLIP encoders and LLMs Liu et al. (2023b; a; 2024); Dai et al. (2023); Bai et al. (2023b); Hong et al. (2023); Wang et al. (2023); Sun et al. (2023a), mixture-of-expert-LLMs Lin et al. (2024), optimized architectures for resource-constrained systems Kim et al. (2023), finetuning with RLHF Sun et al. (2023c); OpenAI (2023), or massive vision encoders Chen et al. (2024). Additionally, we survey commercial, closed-source models like Gemini Pro Vision 1.0 Gemini Team (2023), GPT-4V (Preview) OpenAI (2023), and Qwen-VL Plus/Max Qwen Team (2024) where access is limited to APIs and few details are known. For a detailed list of models, please refer to Appendix A.
4.1 Key Results